Why is the order of evaluation of arithmetic expressions important? Illustrate your answer with an example showing the effect parentheses can have on the way an expression is evaluated.
1
Expert's answer
2015-01-19T05:05:28-0500
Answer:
Arithmetic operations in Visual Basic executes accordingto the common arithmetic rules:
1) Compiler evaluates expression fromleft to right
2) Multiplication and division havebigger priority then addition and substraction
3) Parentheses change evaluation order

Examples:
result =3 + 2 * 2 ' resultis 7
result =(3 + 2) * 2 ' resultis 10
result =3 + 2 * 4 / 2 ' resultis 7, first 2* 4, then 8 / 2, then 3 + 4
result =3 + 2 * 4 / 2 * 2 'result is 11, first -> 2* 4, then 8 / 2, then 4 * 2, then 3 + 8
result = 3 + 2 * 4 / (2 * 2) ' result is 5, first -> 2* 2, then 2 * 4, then 8 / 4, then 3+2

Dim a, b, c, d, e, f, g As Double
a = 8.0
b = 3.0
c = 4.0
d = 2.0
e = 1.0
f = (a - b) + ((c / d) * e) = (8 - 3) + ((4 / 2) * 1) = 7
Parentheses priority still from left to right andcannot be changed.